Use the following information to perform the calculations requested below, assuming that the FIFO costing approach to a process costing system is being used. Beginning work in process inventory:

10,000 units, 100% percent complete for direct materials

65 percent complete as to conversion costs

Ending work in process inventory: 2,000 units, 100 percent complete for direct materials

40 percent complete as to conversion costs

Units started during the month: 9,000

a. Calculate equivalent units for materials

b. Calculate equivalent units for conversion costs.
Answer #1

Unit transferred out = 10000+9000-2000 = 17000

a) Equivalent unit of material = (10000*0%+7000*100%+2000*100%) = 9000

b) Equivalent unit of conversion = (10000*35%+7000+2000*40%) = 11300